# RobMOT: Robust 3D Multi-Object Tracking by Observational Noise and State Estimation Drift Mitigation on LiDAR PointCloud

### v3

@@ 3.1.2 Qualitative Evaluation
Although the quantitative evaluation shows the overall performance of a tracking approach, it is not sufficient to evaluate the method under challenging tracking conditions such as occlusions. Thus, we select two challenging conditions to emphasize the contribution of the proposed approach. The first condition represented in the first two rows in Figure 7 is for a distant car that faces frequent occlusions for about 90 frames. The detector is unified for both methods, ours and CasTrack [1], to measure the contribution of the tracking approach. Our proposed solution maintains tracking for the car with ID $41$ to frame $193$ when the car leaves the field of view of LiDAR, while CasTrack [1] lost the car at frame $180$.
@@ 3.2.1 Trajectory Drift
Metric $D_{t}$ VirConv CasA PointRCNN PV-RCNN Second HOTA Y 83% 80.5% 75.7% 78.1% 76.6% N 80.5% 77.2% 72.1% 73.4% 73.5% Improvement - +2.5% +3.3% +3.6% +4.7% +3.1% MOTA Y 87.2% 86.4% 81.5% 84.2% 83.2% N 82.4% 80.8% 76% 77.6% 78.8% Improvement - +4.8% +5.6% +5.5% +6.6% +4.4% IDSW Y 8 4 11 9 8 N 11 19 40 44 33 Improvement - -3 -15 -29 -35 -25 MT Y 479 509 501 496 492 N 470 498 477 484 483 Improvement - +9 +11 +24 +12 +9
In this experiment, we compare the performance of the Kalman filter without the proposed refinement for noise mitigation, $D_{t}$, and the contribution by including the refined Kalman filter. The experiment is conducted on five different detectors to justify the contribution of the proposed refinement Kalman filter, as demonstrated in Table 3. The comparison focuses on metrics MOTA, HOTA, IDSW, and the number of mostly tracked objects (MT). Since our method is not a learning approach, the evaluation is conducted on the KITTI training and evaluation dataset of $21$ streams in total. The proposed refinement in Kalman filter consistently enhances the tracking performance across the five detectors, as shown in Table 3. The improvement in HOTA reached $4.7\%$ with PV-RCNN [23] detector and $6.6\%$ in MOTA. Moreover, it shows consistent state estimation for trajectories that lead to fewer ID switches (IDSW), reduction by $35$ IDs, and more tracked objects (MT); about $24$ objects increased with the PointRCNN [22] detector.
These results, including the qualitative results represented in this work, confirm the contribution of the proposed Kalman filter refinement that mitigates the noise associated with the detections, enhancing state estimation and tracking performance.
@@ 3.2.2 Track Validity and Ghost Tracks
The second concern raised in this research is ghost tracks. Ghost tracks are a form of consecutive false positive observations that generate illusion trajectories that could mislead MOT methods. This work proposes an online trajectory validation mechanism to identify ghost tracks among legitimate trajectories in the system. We validate the contribution of the proposed mechanism in two steps. First, we demonstrate a qualitative scenario where ghost and legitimate tracks exist and the capability of the mechanism to identify legitimate tracks. Second, we perform a sensitivity analysis by disabling the mechanism and quantifying the improvement in the tracking performance.
@@
Detector Traj. Validity HOTA $\uparrow$ MOTA $\uparrow$ IDSW $\downarrow$ PointRCNN [22] N 69.38% 63.59% 58 PointRCNN [22] Y 75.66% 81.46% 11 Improvement - +6.28% +17.87% -47 PV-RCNN [23] N 73.92% 73.72% 65 PV-RCNN [23] Y 78.13% 84.2% 9 Improvement - +4.21% +10.45% -56 Second [21] N 73.69% 75.78% 39 Second [21] Y 76.64% 83.22% 8 Improvement - +2.95% +7.44% -31 CasA [20] N 78.28% 80.75% 26 CasA [20] Y 80.53% 86.36% 4 Improvement - +2.25% +5.61% -22 VirConv [19] N 80.66% 83.03% 21 VirConv [19] Y 83.01% 87.17% 8 Improvement - +2.35% +4.14% -13
To support the proposed trajectory validity mechanism’s impact on tracking performance, we disable the mechanism and quantitatively evaluate the tracking performance on the training and evaluation KITTI dataset. To demonstrate the generalization of the proposed mechanism, we employ five different detectors and quantify the enhancement in performance across five detectors when the trajectory validity mechanism is enabled. While the trajectory validity mechanism is disabled, we employ threshold on the detection score as proposed and fine-tuned in CasTrack [1] work. The evaluation involves HOTA, MOTA, and IDSW metrics to represent the overall tracking performance. Table 4 shows that enabling the trajectory validity mechanism improves HOTA across all detectors, with a margin reaching $6.28\%$ with PointRCNN [22] detector. Moreover, there is a significant improvement in the multi-object tracking accuracy (MOTA) metric for overall detectors that peak with the PointRCNN detector with up to $17.87\%$ enhancement in MOTA. Since the trajectory validity mechanism identifies legitimate tracks while allowing their observations to pass through the system, the tracking performance becomes more consistent with less misidentification for objects that reflect few IDSW values, as shown in Table 4. It results in a reduction in IDSW reaches $56$ with PV-RCNN [23] detector.
@@ 4 Conclusion
This paper discusses two critical challenges in multi-object tracking (MOT): state estimation deviation and ghost tracks. Our research reveals that the noise associated with detections from the detector affects the accuracy of bounding box localization, leading to errors in estimating the motion parameters of these objects. This problem causes trajectory estimation to drift, particularly when objects encounter occlusion, and has been noted in recent advanced MOT methods. Another type of noise is formed from false positive observations that lead to ghost tracks. While the literature relies on fine-tuned thresholds on the detection score to filter observations contributing to these tracks, it still leaks false observations and blocks observation of legitimate tracks, eventually impacting tracking performance. We addressed this issue with a refined Kalman filter compensating for trajectory drift noise, allowing drift mitigation, and recovering prolonged occluded objects. The work also introduces a novel online trajectory validity mechanism to identify legitimate tracks based on ghost track characteristics in MOT methods.
The proposed solutions are combined in an online MOT framework, RobMOT, that outperforms state-of-the-art trackers in KITTI and Waymo Open Dataset. The framework shows superior tracking performance across various detectors with a margin reaching $4\%$ in HOTA and $6\%$ in MOTA with the top-performing MOT benchmark. The framework qualitatively shows its substantial tracking performance for distant and recovering prolonged occluded objects when SOTA fails. An ablation study has been conducted to demonstrate the improvement in the tracking performance resulting from each component proposed in this work. The study shows that mitigation of the state estimation drift noise enhances HOTA by $4.7\%$ and MOTA by 6.6%. Meanwhile, the proposed trajectory validity mechanism significantly improves the tracking performance, reaching $6.28\%$ in HOTA and 17.87% in MOTA, dramatically reducing the IDSW metric to $56$ IDs. The framework shows its applicability to real-time solutions with CPU constraints and high computational processing.
Appendix A Thresholds and Parameter Values
@@
Detector $D_{t}(\sigma_{x}^{2})$ $D_{t}(\sigma_{y}^{2})$ $\sigma_{nconf}$ $\sigma_{conf}$ $\sigma_{legit}$ $\sigma_{conv}$ VirConv [19] 0.016629 0.005334 0 -1 20 4 CasA [20] 0.030696 0.015416 0 0 25 4 PointRCNN [22] 0.032043 0.009945 0 0 35 4 PV-RCNN [23] 0.034076 0.012463 0.5 0.5 20 4 Second [21] 0.037623 0.013561 -1 -2 10 4

### v4

@@ 4.2.2 Qualitative Evaluation
Although the quantitative evaluation shows the overall performance of a tracking approach, it is not sufficient to evaluate the method under challenging tracking conditions such as occlusions. Thus, two challenging conditions are selected to emphasize the contribution of the proposed approach. The first condition represented in the first two rows in Figure 8 is for a distant car that faces frequent occlusions for about 90 frames. The detector is unified for both methods, ours and CasTrack [1], to measure the contribution of the tracking approach. Our proposed solution maintains tracking for the car with ID $41$ to frame $193$ when the car leaves the field of view of LiDAR, while CasTrack [1] lost the car at frame $180$.
@@
4.2.3 Runtime and Speed Comparison
A third study is conducted to provide some insights into the computational efficiency of the proposed approach. Since the number of incoming detections and detections filtration in the MOT method are the main contributors to the run-time efficiency, this experiment compares the performance of the proposed method (RobMOT) with the recent state-of-the-artwork that achieves state-of-the-art tracking speed (Figure 4 in [1]) across five different detectors. In this study, the method [1, 3] is re-run on the same machine as the proposed approach (RobMOT) with hardware stated in Section 4.1. Figure 9 summarizes the frame per second (FPS) achieved by each method on training and validation KITTI datasets that contains about $7,500$ frames. Our proposed solution significantly reduces computational time compared to method [1, 3] across the five detectors. Our approach achieves an FPS margin that reaches $2000$ FPS, as shown with the VirConv detector [2]. This result reflects the applicability of the proposed method for real-time applications, particularly applications with hardware constraints.
Figure 9: Runtime comparison of processing about 7,500 consecutive frames from KITTI dataset, training and validation splits, across five detectors with benchmark [1, 3].
@@ 4.3.1 Trajectory Drift
Metric $D_{t}$ VirConv CasA PointRCNN PV-RCNN Second HOTA Y 85.1% 82.2% 76.7% 78.6% 77.2% N 82.7% 79% 73.2% 73.8% 73.8% Improvement - +2.4% +3.2% +3.5% +4.8% +3.4% MOTA Y 91.4% 90.3% 86% 86.8% 85.3% N 91% 89.4% 84.8% 85.5% 83.9% Improvement - +0.4% +0.9% +1.2% +1.3% +1.4% IDSW Y 1 0 3 1 5 N 2 5 10 14 10 Improvement - -1 -5 -7 -13 -5 MT Y 159 162 152 153 155 N 159 160 146 151 151 Improvement - 0 +2 +6 +2 +4
In this experiment, the performance of the KF without the proposed refinement for noise mitigation, $D_{t}$, is compared with the proposed refined KF. The experiment is conducted on five different detectors to justify the contribution of the proposed refinement KF, as demonstrated in Table 3. The comparison focuses on metrics MOTA, HOTA, IDSW, and the number of mostly tracked objects (MT). For this ablation study, the noise mitigation term $D_{t}$ is modeled on the KITTI training dataset to measure the generalization of the mitigation term on the unseen dataset, which is the KITTI evaluation dataset in this case. Since this study aims to evaluate the impact of localization and noise mitigation on state estimation, Table 3 compares state prediction for objects (Before refining the state given the new measurement) and the actual location of the objects from the ground truth. Thus, the improvement row in Table 3 reflects enhancement in state prediction due to including the noise mitigation term $D_{t}$.
The proposed refinement in the KF consistently enhances the tracking performance across the five detectors, as shown in Table 3. Notably, the improvement increases as the detection performance decreases. Table 3 shows significant improvement in HOTA reaches $4.8\%$ with PV-RCNN [28] detector. The primary reason for the significant improvement in HOTA compared to MOTA is that HOTA includes localization accuracy, while MOTA considers identity switch between objects (See Section 4.1). This observation shows the contribution of the proposed noise mitigation term $D_{t}$ that leads to enhanced state prediction for objects. The improvement in state prediction allows recovering objects that reach to recover about 13 objects with PV-RCNN [28] detector (See IDSW in Table 3).
These results, including the qualitative results represented in this work, confirm the contribution of the proposed KF refinement that mitigates the noise associated with the detections, enhancing state estimation and tracking performance.
@@ 4.3.2 Track Validity and Ghost Tracks
The second concern raised in this research is ghost tracks. Ghost tracks are a form of consecutive false positive observations that generate illusion trajectories that could mislead MOT methods. This work proposes the first online trajectory validation mechanism to identify ghost tracks among legitimate trajectories in MOT methods. The contribution of the proposed mechanism is validated in two steps. First, a qualitative scenario is demonstrated in which ghost and legitimate tracks exist and the capability of the mechanism to identify legitimate tracks. Second, a sensitivity analysis is performed by disabling the mechanism and quantifying the improvement in the tracking performance.
@@
Detector Traj. Validity HOTA $\uparrow$ MOTA $\uparrow$ IDSW $\downarrow$ PointRCNN [28] N 69.37% 58.54% 16 PointRCNN [28] Y 78% 86.55% 3 Improvement - +8.63% +28.01% -13 PV-RCNN [29] N 74.58% 71.02% 35 PV-RCNN [29] Y 79.92% 86.93% 1 Improvement - +5.34% +15.91% -34 Second [11] N 69.80% 56.26% 30 Second [11] Y 78.73% 85.73% 7 Improvement - +8.93% +29.47% -23 CasA [4] N 81.72% 86.13% 10 CasA [4] Y 83.61% 90.58% 0 Improvement - +1.89% +4.45% -10 VirConv [2] N 85.87% 90.88% 10 VirConv [2] Y 86.31% 91.53% 1 Improvement - +0.44% +0.65% -9
To support the proposed trajectory validity mechanism’s impact on tracking performance, the mechanism is turned off and quantitatively evaluates the tracking performance on the KITTI evaluation dataset after fine-tuning the mechanism on the training dataset. Five different detectors are employed to demonstrate the generalization of the proposed mechanism and the performance enhancement across five detectors is quantified when the trajectory validity mechanism is enabled. In this study, the trajectory validity mechanism threshold $\sigma_{legit}$ is initially finetuned on the KITTI training dataset, and its performance is evaluated on the KITTI evaluation dataset. The evaluation is conducted by turning off the proposed mechanism and employing threshold on the detection score as proposed and finetuned in CasTrack [1] work. Next, the impact of the proposed mechanism can be observed when enabled.
The evaluation involves HOTA, MOTA, and IDSW metrics to represent the overall tracking performance. Table 4 shows that enabling the trajectory validity mechanism improves HOTA across all detectors, with a margin reaching $8.94\%$ with Second [11] detector. Moreover, there is a significant improvement in the MOTA metric for overall detectors that peak with the PointRCNN detector with up to $28.01\%$ enhancement in MOTA. Since the trajectory validity mechanism identifies legitimate tracks while allowing their observations to pass through the system, the tracking performance becomes more consistent with less misidentification for objects that reflect few IDSW values, as shown in Table 4. It results in a reduction in IDSW reaches $34$ with PV-RCNN [29] detector.
@@
5 Limitations and Future Work
This work observes trajectory drift for occluded objects, which fails MOT approaches to recover these objects after the end of the occlusion. This research shows that this drift results from noise from the alimitation and futuressigned deep-learning detector that impacts the localization of the detected objects. Accumulating this noise shapes a Gaussian distribution that facilitates the integration with the KF. The proposed term mitigates the need for drift-enhancing state estimation for objects, which leads to recovering objects under extended occlusion scenarios. Nevertheless, this term assumes the impact of this noise on objects is almost equal, even though this noise can differ as per the view perspective of the object, including the number of observations for that object. Although the error that comes from that assumption is partially filtered during state estimation through the KF, dismissing that assumption by a dynamic formulation for this noise can allow further enhancement in state estimation.
Another contribution introduced in this work is to overcome the formulation of ghost tracks in MOT by replacing the traditional threshold-based filter approach with a trajectory validity mechanism that provides real-time distinction between legitimate and ghost tracks in MOT approaches. As shown in the ablation study in Section 4.3.2, the mechanism significantly enhances tracking performance in the MOTA metric, which indicates a reduction in false-positive tracks in the method. Moreover, this reduction in false tracks prevents further mismatching for objects’ identities showing less IDSW, as presented in Table 1. Nevertheless, the mechanism requires multiple observations to validate the legitimacy of a track, which means this constraint may exclude legitimate tracks for objects with too fewer observations, i.e., only one or two observations for the object. Our observation of this limitation concludes that most of these objects are already leaving the scene of the AV; keeping their track in the MOT method does not provide important information for the AV. Nevertheless, this observation triggers the question, "Which objects should be tracked?". Current research in MOT assumes that all objects must be tracked; however, this is not always true for real-time AV environments. For instance, tracking information for cars driving on a road different than the AV is not always beneficial for the AV to navigate safely, but overloading the MOT method instead can eventually harm the tracking performance of other necessary tracks.
Lastly, this work uses Kalman with a constant acceleration motion model for state estimation; however, this assumption does not apply to all objects. For instance, vehicles regularly change their motion according to traffic. This assumption also impacts the precision of state estimation for objects in the KF.
@@ 6 Conclusion
This paper addresses two key challenges in multi-object tracking (MOT): state estimation deviation and ghost tracks. This research finds that spatial localization noise from deep learning detectors impacts bounding box localization accuracy, leading to imprecise motion parameter estimation and trajectory drift. This issue is observed in multiple scenarios, including the recent benchmarks. Moreover, this research discusses the limitation of the current approach followed in the literature, threshold-based filtration, to filter out ghost tracks (false positives) and its limitation of leaking ghost tracks to the MOT approach while blocking legitimate ones. This work tackles these limitations by introducing a refined Kalman filter that compensates for the spatial localization noise, enabling trajectory drift mitigation and recovery of prolonged occluded objects. In addition, The work introduces a novel online trajectory validity mechanism to identify legitimate tracks based on ghost track characteristics. Finally, the work proposes an online MOT framework, RobMOT, which outperforms state-of-the-art trackers on the KITTI and Waymo Open Dataset (WOD). RobMOT shows superior tracking performance across various detectors, with HOTA improvements up to 3.92% and 8.7% on the KITTI testing and validation dataset. Compared to SOTA methods, the framework excels in tracking distant and recovering occluded objects with a distance exceeding 50 meters, followed by a 1.27% improvement in MOTA on the WOD testing split. The framework shows computational efficiency that reflects its applicability for real-time applications.
@@
Detector $D_{t}(\sigma_{x}^{2})$ $D_{t}(\sigma_{y}^{2})$ $\sigma_{nconf}$ $\sigma_{conf}$ $\sigma_{legit}$ $\sigma$ VirConv [2] 0.017221 0.005901 0 -1 20 4 CasA [4] 0.034966 0.019720 0 0 25 3 PointRCNN [28] 0.030874 0.009379 0 0 35 4 PV-RCNN [29] 0.036383 0.013067 0.5 0.5 20 2 Second [11] 0.039156 0.014357 -1 -2 10 3
The trajectory is discarded from the cache for all detectors when the estimation uncertainty $diag(P_{t})_{x}$ or $diag(P_{t})_{y}$ exceeds $\sigma_{est\_certainity}=4$ meters.